# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ Appendix B Proofs in Section 3
Theorem 3.20 (Gradient Formula in the Field). Given a local objective operator ${\bm{\Gamma}}_{x}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, the gradient formula for ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ w.r.t. local generator ${{\mathbf{G}}}(x)$ is
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)_{{\omega}}^{{\omega}^{\prime}}}={\bm{\Gamma}}_{x}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\hat{\bar{\bm{\varphi}}}}.$ (B.73)
@@
  • ${{\mathbf{P}}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$  for every acting entity $x^{\prime}$ connecting to acting entity $x$;
  • ${\bm{\Gamma}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bar{\gamma}}$  for every environmental entity $x^{\prime}$ connecting to acting entity $x$;
@@
$\forall x\in{\mathcal{X}}_{a}:{\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bar{\gamma}}.$ (B.81)
@@
and ${\bm{\Gamma}}_{x_{0}x_{i}}$ is given by the environmental entity. Multiplying both sides by ${\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$, we then apply the first two conditions as stated in the proposition:
${\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=\sum_{j=0}^{n}{\Lambda}^{x_{j}}_{x_{i}}{\bm{\Gamma}}_{x_{j}x_{i}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\Lambda}^{x_{0}}_{x_{i}}{\bar{\gamma}}+\sum_{j=1}^{n}{\Lambda}^{x_{j}}_{x_{i}}{\bm{\Gamma}}_{x_{j}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (B.83)
Denote $\Theta\in{\mathbb{R}}^{n\times n}$ as a matrix where its elements given by $\Theta_{ij}:=\delta_{i}^{j}-{\Lambda}_{x_{i}}^{x_{j}}$. Moreover, denote $\bm{a},\bm{b}\in{\mathbb{R}}^{n}$ as two $n$-dimensional vectors where $a_{i}:={\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}\in{\mathbb{R}}$ and $b_{i}={\Lambda}^{x_{0}}_{x_{i}}{\bar{\gamma}}$. Then, the above equation can be formulated into
@@
${\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bar{\gamma}}.$ (B.85)
@@
    ${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta{{\mathbf{P}}}$ satisfies $\delta{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$.
@@
We are going to use two facts. The first fact is ${{\mathbf{G}}}={\mathbf{\Pi}}{\widetilde{\mathbf{G}}}$, and the second is ${{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$.
@@
${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (B.89)
@@
Note that, with the ergodicity, ${\mathbf{\Phi}}$ maps everything to ${\hat{\bar{\bm{\varphi}}}}$, i.e., ${\mathbf{\Phi}}:{\Delta\mathcal{H}}\to 0$. In addition, since ${\mathbf{\Phi}}$ and ${{\mathbf{S}}}$ are all in the form of $e^{{{\mathbf{G}}}t}$, they commute. Therefore, ${\mathbf{\Phi}}{{\mathbf{S}}}{\mathbf{\Pi}}={{\mathbf{S}}}{\mathbf{\Phi}}{\mathbf{\Pi}}=0$. This proves
@@
$\delta{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}){\mathbf{\Pi}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}){{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0.$ (B.95)
@@
    ${{\mathbf{P}}}[{{\mathbf{Q}}}]{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}])$ satisfies $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}]){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$.
@@
The proof is very similar to that of Proposition 3.28. We are going to use two facts. The first fact is ${{\mathbf{G}}}={\mathbf{\Pi}}{\widetilde{\mathbf{G}}}$, and the second is ${{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$.
@@
${{\mathbf{P}}}[{{\mathbf{Q}}}]{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}{\mathbf{\Pi}}){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}{{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (B.96)
@@
$\delta({{\mathbf{P}}}[{{\mathbf{Q}}}]){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}){\mathbf{\Pi}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}){{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0.$ (B.103)
@@
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}_{x}{{\mathbf{P}}}{{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}{\hat{\bar{\bm{\varphi}}}}.$ (B.104)
@@
The key ingredient in proving the theorem is to show that $\delta{\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$ for propagator ${{\mathbf{P}}}_{x^{\prime}x}={{\mathbf{P}}}[{{\mathbf{Q}}}_{x^{\prime}x}]$. We may view the collection of all environmental entities as a single entity by relabeling configurations sets. Then, the “boundary conditions” of the objective propagation equation are given by the objective signals from the environmental entity, which stays unchanged. We may label the entities by $x_{0},x_{1},\dots,x_{n}$ where $x_{0}$ denotes the environmental entity, and the rest $n$ entities are the acting entities. Thus, the objective propagation equation (Definition 3.24) can be written as
@@
Then, The above equation can be very much simplified by multiplying ${\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ on both sides and applying Proposition 3.31.
$\delta{\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ $=\sum_{j=1}^{n}{\Lambda}_{x_{i}}^{x_{j}}\left({\bm{\Gamma}}_{x_{j}}\delta({{\mathbf{P}}}[{{\mathbf{Q}}}_{x_{j}x_{i}}]){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}+\delta{\bm{\Gamma}}_{x_{j}}{{\mathbf{P}}}[{{\mathbf{Q}}}_{x_{j}x_{i}}]{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}\right)$ (B.107)
$=\sum_{j=1}^{n}{\Lambda}_{x_{i}}^{x_{j}}\delta{\bm{\Gamma}}_{x_{j}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (B.108)
Denote $\Theta\in{\mathbb{R}}^{n\times n}$ as a matrix where its elements given by $\Theta_{ij}:=\delta_{i}^{j}-{\Lambda}_{x_{i}}^{x_{j}}$. Moreover, denote $\bm{a}\in{\mathbb{R}}^{n}$ as a vector where $a_{i}:=\delta{\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}\in{\mathbb{R}}$. Then, the above equation can be formulated into
@@
$\delta{\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0.$ (B.110)
@@
$\delta{\bar{\gamma}}=\delta({\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}})={\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}})+\delta{\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}).$ (B.111)

### v2

@@ Appendix B Proofs in Section 3
Theorem 3.20 (Gradient Formula in the Field). Given a local objective operator ${\bm{\Gamma}}_{x}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, the gradient formula for ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ w.r.t. local generator ${{\mathbf{G}}}(x)$ is
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)_{{\omega}}^{{\omega}^{\prime}}}={\bm{\Gamma}}_{x}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\lvert\bar{\varphi}\rangle}.$ (B.73)
@@
  • ${{\mathbf{P}}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ for every acting entity $x^{\prime}$ connecting to acting entity $x$;
  • ${\bm{\Gamma}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bar{\gamma}}$ for every environmental entity $x^{\prime}$ connecting to acting entity $x$;
@@
$\forall x\in{\mathcal{X}}_{a}:{\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bar{\gamma}}.$ (B.81)
@@
and ${\bm{\Gamma}}_{x_{0}x_{i}}$ is given by the environmental entity. Multiplying both sides by ${\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$, we then apply the first two conditions as stated in the proposition:
${\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=\sum_{j=0}^{n}{\Lambda}^{x_{j}}_{x_{i}}{\bm{\Gamma}}_{x_{j}x_{i}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\Lambda}^{x_{0}}_{x_{i}}{\bar{\gamma}}+\sum_{j=1}^{n}{\Lambda}^{x_{j}}_{x_{i}}{\bm{\Gamma}}_{x_{j}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (B.83)
Denote $\Theta\in{\mathbb{R}}^{n\times n}$ as a matrix where its elements given by $\Theta_{ij}:=\delta_{i}^{j}-{\Lambda}_{x_{i}}^{x_{j}}$. Moreover, denote $\bm{a},\bm{b}\in{\mathbb{R}}^{n}$ as two $n$-dimensional vectors where $a_{i}:={\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}\in{\mathbb{R}}$ and $b_{i}={\Lambda}^{x_{0}}_{x_{i}}{\bar{\gamma}}$. Then, the above equation can be formulated into
@@
${\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bar{\gamma}}.$ (B.85)
@@
    ${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta{{\mathbf{P}}}$ satisfies $\delta{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$.
@@
We are going to use two facts. The first fact is ${{\mathbf{G}}}={\mathbf{\Pi}}{\widetilde{\mathbf{G}}}$, and the second is ${{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$.
@@
${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (B.89)
@@
Note that, with the ergodicity, $\mathbf{\Phi}$} maps everything to $\lvert\bar{\varphi}\rangle$}, i.e., ${\mathbf{\Phi}}:{\Delta\mathcal{H}}\to 0$. In addition, since $\mathbf{\Phi}$} and $\mathbf{S}$}} are all in the form of $e^{{{\mathbf{G}}}t}$, they commute. Therefore, ${\mathbf{\Phi}}{{\mathbf{S}}}{\mathbf{\Pi}}={{\mathbf{S}}}{\mathbf{\Phi}}{\mathbf{\Pi}}=0$. This proves
@@
$\delta{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}){\mathbf{\Pi}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}){{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0.$ (B.95)
@@
    ${{\mathbf{P}}}[{{\mathbf{Q}}}]{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}])$ satisfies $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}]){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$.
@@
The proof is very similar to that of Proposition 3.28. We are going to use two facts. The first fact is ${{\mathbf{G}}}={\mathbf{\Pi}}{\widetilde{\mathbf{G}}}$, and the second is ${{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$.
@@
${{\mathbf{P}}}[{{\mathbf{Q}}}]{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}{\mathbf{\Pi}}){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}{{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (B.96)
@@
$\delta({{\mathbf{P}}}[{{\mathbf{Q}}}]){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}){\mathbf{\Pi}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=\delta({\widetilde{\mathbf{G}}}{{\mathbf{S}}}{{\mathbf{Q}}}){{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0.$ (B.103)
@@
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}_{x}{{\mathbf{P}}}{{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}{\lvert\bar{\varphi}\rangle}.$ (B.104)
@@
The key ingredient in proving the theorem is to show that $\delta{\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$ for propagator ${{\mathbf{P}}}_{x^{\prime}x}={{\mathbf{P}}}[{{\mathbf{Q}}}_{x^{\prime}x}]$. We may view the collection of all environmental entities as a single entity by relabeling configurations sets. Then, the “boundary conditions” of the objective propagation equation are given by the objective signals from the environmental entity, which stays unchanged. We may label the entities by $x_{0},x_{1},\dots,x_{n}$ where $x_{0}$ denotes the environmental entity, and the rest $n$ entities are the acting entities. Thus, the objective propagation equation (Definition 3.24) can be written as
@@
Then, The above equation can be very much simplified by multiplying ${\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ on both sides and applying Proposition 3.31.
$\delta{\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ $=\sum_{j=1}^{n}{\Lambda}_{x_{i}}^{x_{j}}\left({\bm{\Gamma}}_{x_{j}}\delta({{\mathbf{P}}}[{{\mathbf{Q}}}_{x_{j}x_{i}}]){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}+\delta{\bm{\Gamma}}_{x_{j}}{{\mathbf{P}}}[{{\mathbf{Q}}}_{x_{j}x_{i}}]{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}\right)$ (B.107)
$=\sum_{j=1}^{n}{\Lambda}_{x_{i}}^{x_{j}}\delta{\bm{\Gamma}}_{x_{j}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (B.108)
Denote $\Theta\in{\mathbb{R}}^{n\times n}$ as a matrix where its elements given by $\Theta_{ij}:=\delta_{i}^{j}-{\Lambda}_{x_{i}}^{x_{j}}$. Moreover, denote $\bm{a}\in{\mathbb{R}}^{n}$ as a vector where $a_{i}:=\delta{\bm{\Gamma}}_{x_{i}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}\in{\mathbb{R}}$. Then, the above equation can be formulated into
@@
$\delta{\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0.$ (B.110)
@@
$\delta{\bar{\gamma}}=\delta({\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle})={\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle})+\delta{\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}).$ (B.111)